### v3

@@ 4.3.2 VaR and ES forecasts
At the 5% level, the unconditional test shows that SV-GH-ST is significantly outperformed by SV-T, SV-AZ-SN, SV-AZ-ST, and all RSV models. In addition, RSV-GH-ST outperforms SV-N, and RSV-AZ-ST significantly outperforms RSV-AZ-SN. Among EGARCH models, EG-T again significantly outperforms EG-N. In the REGARCH group, REG-T outperforms REG-N, and both REGARCH variants outperform SV-GH-ST.
@@
For the conditional GW test at the 5% level, SV-GH-ST is significantly outperformed by RSV-AZ-SN, RSV-AZ-ST, RSV-FS-SN, RSV-FS-ST, and RSV-GH-ST. RSV-T performs significantly worse than RSV-GH-ST. EG-N is significantly outperformed by EG-T, and REG-N by REG-T.
@@
In summary, the results highlight the superior performance of the RSV models—especially RSV-AZ-ST, RSV-FS-ST, and RSV-GH-ST—compared to SV, EGARCH, and REGARCH models, particularly under stringent risk levels. The inclusion of RV consistently improves the accuracy of tail risk forecasts. Moreover, as the FZ0 loss function jointly evaluates VaR and ES by emphasizing tail behavior, these results underscore the importance of capturing skewness and heavy tails in return distributions.
Table 7: Summary of GW tests on FZ0 loss of VaR and ES forecasts for DJIA
$\alpha=1\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $94$ $99$ $99$ $100$ $99$ $6$ $1$ $1$ $9$ $54$ $99$ $12$ $99$ $1$ $23$ $1$ $1$
SV-T $-0.4$ $99$ $100$ $35$ $100$ $7$ 00 00 00 $1$ $14$ $1$ $99$ $2$ $2$ 00 00
SV-AZ-SN $-2.2$ $-1.6$ $4$ $6$ 00 $2$ 00 00 00 $1$ 00 00 $1$ $1$ $1$ 00 00
SV-AZ-ST $-0.9$ $-1.0$ $1.4$ $47$ $59$ $4$ 00 00 00 00 $1$ 00 $12$ $2$ $1$ 00 00
SV-FS-SN $-1.5$ $-0.8$ $0.2$ $-0.6$ $11$ $2$ 00 00 $5$ $16$ 00 $4$ $49$ 00 00 00 00
SV-FS-ST $-1.4$ $-0.9$ $0.9$ $-0.7$ $0.3$ $2$ 00 00 00 00 00 00 $4$ $1$ $1$ 00 00
SV-GH-ST $0.5$ $0.8$ $2.0$ $1.5$ $1.5$ $1.9$ $2$ $2$ $4$ $89$ $96$ $57$ $97$ $2$ $71$ 00 00
RSV-N $1.1$ $1.4$ $1.6$ $1.5$ $1.4$ $1.4$ $0.6$ $100$ $100$ $100$ $100$ $100$ $100$ $2$ $99$ $2$ $2$
RSV-T $0.8$ $1.1$ $1.4$ $1.2$ $1.2$ $1.2$ $0.3$ $-2.2$ $100$ $99$ $100$ $100$ $100$ $2$ $99$ 00 00
RSV-AZ-SN $0.4$ $0.8$ $1.2$ $0.9$ $1.0$ $1.0$ $0.0$ $-2.3$ $-1.6$ $100$ $97$ $97$ $100$ $2$ $99$ 00 00
RSV-AZ-ST $0.1$ $0.3$ $1.0$ $0.6$ $0.8$ $0.7$ $-0.3$ $-2.3$ $-2.0$ $-1.7$ $90$ 00 $100$ $2$ $62$ 00 00
RSV-FS-SN $-0.2$ $0.0$ $0.9$ $0.3$ $0.7$ $0.6$ $-0.6$ $-2.1$ $-1.8$ $-1.2$ $-0.6$ $1$ $100$ $2$ $6$ 00 00
RSV-FS-ST $0.2$ $0.5$ $1.1$ $0.8$ $0.9$ $0.8$ $-0.2$ $-2.4$ $-2.1$ $-1.2$ $0.9$ $1.1$ $100$ $2$ $98$ 00 00
RSV-GH-ST $-0.6$ $-0.4$ $0.6$ $0.0$ $0.4$ $0.3$ $-0.9$ $-2.2$ $-2.0$ $-1.8$ $-1.7$ $-0.8$ $-1.8$ $2$ $1$ 00 00
EG-N $1.5$ $1.8$ $1.9$ $1.8$ $1.6$ $1.7$ $0.9$ $0.4$ $0.8$ $1.1$ $1.3$ $1.5$ $1.3$ $1.7$ $100$ $1$ $1$
EG-T $0.1$ $0.4$ $1.2$ $0.7$ $0.9$ $0.9$ $-0.3$ $-1.0$ $-0.6$ $-0.3$ $0.0$ $0.3$ $-0.1$ $0.5$ $-2.1$ 00 00
REG-N $1.4$ $1.7$ $1.8$ $1.7$ $1.5$ $1.6$ $0.9$ $1.3$ $1.8$ $2.2$ $2.2$ $2.0$ $2.1$ $2.2$ $0.2$ $1.4$ $98$
REG-T $1.4$ $1.6$ $1.7$ $1.6$ $1.5$ $1.5$ $0.9$ $1.3$ $1.8$ $2.1$ $2.2$ $1.9$ $2.1$ $2.2$ $0.2$ $1.4$ $-1.8$
$\alpha=5\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $96$ $99$ $98$ $95$ $92$ $1$ $97$ $97$ $98$ $98$ $99$ $98$ $98$ $2$ $80$ $95$ $96$
SV-T $-0.6$ $7$ $5$ $7$ $5$ 00 $99$ $99$ $99$ $99$ $99$ $99$ $99$ $2$ $25$ $98$ $98$
SV-AZ-SN $-0.8$ $0.2$ $86$ $4$ 00 00 $97$ $98$ $98$ $98$ $100$ $99$ $99$ $2$ $38$ $96$ $96$
SV-AZ-ST $-0.6$ $0.2$ $-0.2$ $8$ $4$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $3$ $32$ $97$ $97$
SV-FS-SN $-0.1$ $0.5$ $0.8$ $0.5$ $34$ $1$ $96$ $97$ $97$ $98$ $98$ $98$ $98$ 00 $63$ $95$ $96$
SV-FS-ST $-0.1$ $0.6$ $0.9$ $0.7$ $0.1$ 00 $98$ $98$ $98$ $98$ $100$ $100$ $98$ $3$ $53$ $96$ $96$
SV-GH-ST $1.7$ $2.0$ $2.0$ $2.0$ $1.7$ $1.8$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $98$ $98$ $100$ $100$
RSV-N $-1.0$ $-0.6$ $-0.7$ $-0.7$ $-0.9$ $-0.9$ $-2.1$ $99$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $11$ $14$
RSV-T $-1.3$ $-0.9$ $-1.0$ $-0.9$ $-1.1$ $-1.1$ $-2.3$ $-1.7$ $96$ $96$ $82$ $100$ $98$ $1$ $1$ $5$ $6$
RSV-AZ-SN $-1.6$ $-1.3$ $-1.3$ $-1.3$ $-1.4$ $-1.4$ $-2.4$ $-1.4$ $-0.8$ $100$ $73$ $80$ $100$ $2$ $1$ $2$ $3$
RSV-AZ-ST $-1.9$ $-1.6$ $-1.6$ $-1.7$ $-1.7$ $-1.7$ $-2.6$ $-1.8$ $-1.5$ $-2.2$ $1$ $12$ $15$ $2$ $2$ 00 $1$
RSV-FS-SN $-1.7$ $-1.2$ $-1.3$ $-1.3$ $-1.5$ $-1.4$ $-2.4$ $-1.5$ $-0.7$ $0.0$ $1.1$ $100$ $99$ $1$ $2$ $8$ $9$
RSV-FS-ST $-1.8$ $-1.4$ $-1.5$ $-1.5$ $-1.6$ $-1.6$ $-2.5$ $-1.8$ $-1.2$ $-0.7$ $0.8$ $-0.8$ $100$ $2$ $2$ $5$ $7$
RSV-GH-ST $-2.0$ $-1.8$ $-1.8$ $-1.8$ $-1.9$ $-1.9$ $-2.7$ $-1.4$ $-1.1$ $-1.2$ $0.1$ $-0.9$ $-0.6$ $2$ $2$ $1$ $1$
EG-N $0.8$ $1.5$ $1.1$ $1.3$ $0.8$ $0.8$ $-1.1$ $1.4$ $1.6$ $1.9$ $2.1$ $1.9$ $2.0$ $2.2$ $99$ $99$ $99$
EG-T $-0.4$ $0.2$ $0.0$ $0.1$ $-0.4$ $-0.4$ $-1.7$ $0.6$ $0.8$ $1.0$ $1.2$ $1.0$ $1.1$ $1.3$ $-2.3$ $99$ $100$
REG-N $-0.8$ $-0.5$ $-0.5$ $-0.5$ $-0.7$ $-0.7$ $-2.0$ $0.5$ $1.2$ $1.7$ $2.2$ $1.2$ $1.6$ $1.8$ $-1.4$ $-0.4$ $97$
REG-T $-0.8$ $-0.5$ $-0.6$ $-0.6$ $-0.7$ $-0.7$ $-2.0$ $0.3$ $1.0$ $1.5$ $2.1$ $1.1$ $1.5$ $1.7$ $-1.5$ $-0.5$ $-2.5$
  • Notes: See Table 4 for additional details.
Table 8 summarizes the GW test results for the N225, which are broadly consistent with those for the DJIA. At the 1% level, the unconditional test shows that several RSV models significantly outperform SV models, particularly SV-GH-ST, while REGARCH models exhibit mixed performance relative to RSV models. The conditional GW test confirms that RSV models consistently outperform both SV-GH-ST and REGARCH models.
At the 5% level, the unconditional test results indicate that SV-GH-ST is significantly outperformed by all SV models and all RSV models. Although REGARCH models perform better than SV-GH-ST, they are still outperformed by RSV models. The conditional test further supports the dominance of RSV models over SV, EGARCH, and REGARCH variants.
Taken together, the N225 results reinforce the conclusions drawn from the DJIA, providing additional evidence that RSV models—especially RSV-AZ-ST, RSV-FS-ST, and RSV-GH-ST—consistently deliver superior VaR and ES forecasts across varying levels of tail risk.
Table 8: Summary of GW tests on FZ0 loss of VaR and ES forecasts for N225
$\alpha=1\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $1$ $1$ 00 $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ 00 $100$ $100$
SV-T $0.7$ $100$ $99$ $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ 00 $100$ $100$
SV-AZ-SN $0.1$ $-0.5$ $1$ $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $100$ $99$
SV-AZ-ST $0.1$ $-0.4$ $0.0$ $99$ $99$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $100$ $99$
SV-FS-SN $-0.1$ $-0.5$ $-0.3$ $-0.3$ $100$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $100$ $99$
SV-FS-ST $-0.3$ $-0.8$ $-0.5$ $-0.8$ $-0.4$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $1$ $77$ $99$
SV-GH-ST $1.1$ $1.1$ $1.3$ $1.8$ $1.9$ $2.3$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $1$ $94$ $99$ $99$
RSV-N $-1.0$ $-1.1$ $-0.9$ $-0.8$ $-0.7$ $-0.5$ $-1.5$ $99$ 00 $99$ 00 $99$ $99$ 00 00 00 $99$
RSV-T $-1.4$ $-1.5$ $-1.4$ $-1.2$ $-0.9$ $-0.7$ $-1.8$ $-0.5$ 00 $100$ 00 $100$ $100$ 00 00 $1$ $99$
RSV-AZ-SN $-0.8$ $-1.0$ $-0.8$ $-0.7$ $-0.6$ $-0.4$ $-1.4$ $1.0$ $0.7$ $99$ $99$ $99$ $99$ 00 00 $1$ $99$
RSV-AZ-ST $-1.7$ $-2.0$ $-1.7$ $-1.6$ $-1.2$ $-1.0$ $-2.2$ $-0.9$ $-0.5$ $-1.2$ 00 $1$ 00 $1$ 00 $1$ 00
RSV-FS-SN $-1.2$ $-1.4$ $-1.1$ $-1.0$ $-0.8$ $-0.5$ $-1.7$ $0.3$ $1.1$ $0.0$ $1.9$ $100$ $99$ $1$ 00 $2$ $99$
RSV-FS-ST $-1.5$ $-1.7$ $-1.6$ $-1.4$ $-1.1$ $-0.9$ $-2.1$ $-0.8$ $-0.4$ $-1.1$ $0.3$ $-1.7$ $99$ 00 00 $1$ $72$
RSV-GH-ST $-1.6$ $-1.9$ $-1.7$ $-1.6$ $-1.2$ $-1.0$ $-2.2$ $-0.7$ $-0.4$ $-0.9$ $0.3$ $-1.4$ $-0.2$ $1$ 00 $1$ $3$
EG-N $1.0$ $0.8$ $0.8$ $0.8$ $0.7$ $0.9$ $0.1$ $1.7$ $1.8$ $1.6$ $1.8$ $1.6$ $1.7$ $1.7$ $99$ $100$ $99$
EG-T $0.9$ $0.8$ $0.8$ $0.8$ $0.7$ $1.2$ $-0.7$ $1.4$ $1.9$ $1.3$ $2.2$ $1.7$ $2.0$ $2.1$ $-0.7$ $100$ $100$
REG-N $-0.3$ $-0.5$ $-0.3$ $-0.3$ $-0.2$ $0.0$ $-1.0$ $2.2$ $1.5$ $1.8$ $1.7$ $1.0$ $1.7$ $1.5$ $-1.4$ $-0.8$ $99$
REG-T $-1.7$ $-2.2$ $-2.0$ $-2.1$ $-1.6$ $-1.2$ $-2.7$ $-0.4$ $-0.2$ $-0.6$ $0.2$ $-0.9$ $0.0$ $0.1$ $-1.6$ $-2.4$ $-1.0$
$\alpha=5\%$
SV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N T REG-N T
SV-N $99$ $5$ $53$ $97$ $100$ $1$ $99$ $100$ $99$ $100$ $100$ $99$ $99$ $1$ $2$ $98$ $100$
SV-T $-2.0$ 00 $1$ $1$ $3$ 00 $100$ $100$ $99$ $100$ $99$ $100$ $98$ $1$ $2$ $15$ $99$
SV-AZ-SN $0.2$ $3.2$ $97$ $97$ $99$ $1$ $98$ $99$ $98$ $98$ $100$ $98$ $98$ $2$ $2$ $97$ $98$
SV-AZ-ST $-0.1$ $3.2$ $-0.3$ $99$ $86$ 00 $99$ $100$ $99$ $100$ $100$ $99$ $99$ $1$ $2$ $97$ $100$
SV-FS-SN $-0.6$ $1.4$ $-1.1$ $-0.7$ $99$ $1$ $97$ $99$ $97$ $97$ $99$ $97$ $97$ $1$ $1$ $96$ $97$
SV-FS-ST $-0.5$ $1.3$ $-0.8$ $-0.8$ $-0.1$ $1$ $97$ $97$ $97$ $97$ $97$ $97$ $97$ $1$ $6$ $94$ $96$
SV-GH-ST $2.2$ $3.0$ $2.2$ $2.4$ $2.4$ $2.5$ $99$ $99$ $99$ $99$ $99$ $99$ $99$ $96$ $98$ $99$ $99$
RSV-N $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.6$ $99$ $3$ $100$ $15$ $3$ 00 00 00 00 00
RSV-T $-1.3$ $-0.5$ $-1.3$ $-1.2$ $-1.0$ $-0.8$ $-2.7$ $-0.6$ $3$ $98$ 00 $2$ $1$ 00 00 $1$ 00
RSV-AZ-SN $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.6$ $0.1$ $0.7$ $99$ $1$ $2$ 00 00 00 00 00
RSV-AZ-ST $-1.4$ $-0.6$ $-1.3$ $-1.2$ $-1.0$ $-0.9$ $-2.8$ $-0.7$ $-0.3$ $-0.8$ 00 $1$ 00 00 00 00 00
RSV-FS-SN $-1.1$ $-0.3$ $-1.0$ $-0.9$ $-0.7$ $-0.6$ $-2.6$ $0.8$ $1.1$ $0.8$ $1.1$ $99$ $100$ 00 00 $2$ 00
RSV-FS-ST $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.7$ $0.3$ $1.1$ $0.2$ $1.0$ $-0.4$ $38$ 00 00 00 00
RSV-GH-ST $-1.2$ $-0.4$ $-1.1$ $-1.0$ $-0.8$ $-0.7$ $-2.7$ $0.4$ $1.0$ $0.3$ $1.9$ $-0.3$ $0.1$ 00 00 00 $1$
EG-N $1.2$ $2.2$ $1.1$ $1.2$ $1.3$ $1.4$ $-1.3$ $2.0$ $2.2$ $2.1$ $2.2$ $2.0$ $2.1$ $2.1$ $100$ $100$ $100$
EG-T $0.6$ $2.0$ $0.6$ $0.8$ $0.9$ $1.1$ $-2.0$ $1.5$ $1.7$ $1.5$ $1.7$ $1.4$ $1.5$ $1.6$ $-1.5$ $100$ $100$
REG-N $-0.7$ $0.1$ $-0.7$ $-0.6$ $-0.4$ $-0.3$ $-2.4$ $2.0$ $2.7$ $2.1$ $3.0$ $1.3$ $2.1$ $2.6$ $-1.8$ $-1.2$ $98$
REG-T $-0.9$ $-0.1$ $-0.9$ $-0.8$ $-0.6$ $-0.5$ $-2.6$ $1.2$ $1.9$ $1.3$ $2.4$ $0.8$ $1.4$ $1.7$ $-1.8$ $-1.4$ $-0.8$
  • Notes: See Table 4 for additional details.
@@
To explore differences among RSV models, we examine the distributional properties of their one-day-ahead return forecasts. Table 9 reports the mean and standard deviation of skewness, kurtosis, and several lower percentiles (0.1%, 1%, 5%, and 10%) based on the posterior predictive distributions. Figures 12 and 13 present the histograms of these statistics across the forecast horizon for the DJIA and N225, respectively.
@@
4.3.3 Summary
Table 10 presents the rankings of average losses for volatility, VaR, and ES forecasts. The average ranking in the final column highlights the overall superiority of the RSV models, with RSV-AZ-ST achieving the best performance.
Our findings demonstrate that RV is a valuable predictor for forecasting not only volatility but also tail risk measures such as VaR and ES. Across all forecast categories, RSV models consistently outperform REGARCH models. In particular, incorporating a skewed $t$-distribution into the RSV framework appears to further enhance forecast accuracy.
@@ 4.3.3 Summary
SV-N 17 17 17 15   8 16 13 12 13 14 12 14 $14.00$
SV-T 14 14 14 14   6 10 18 18 16 15 15   9 $13.58$
SV-AZ-SN 18 18 18 18   1 13 15 14 14 16 13 15 $14.42$
SV-AZ-ST 15 15 15 16   4 11 17 16 17 17 14 13 $14.17$
SV-FS-SN 11   2   2   2   2 14 10 10 10 11 11 12 $8.08$
SV-FS-ST 10   1   1   1   3 15 12 13 12 12 10 11 $8.42$
SV-GH-ST 16 16 16 17 12 18 16 15 18 18 17 18 $16.42$
RSV-N   6 12 10 11 15   7   3   3   3   5   6   3 $7.00$
RSV-T   3   7   5   5 14   6   5   5   5   3   5   2 $5.42$
RSV-AZ-SN   4 10   8   9 13   4   4   4   4   4   8   4 $6.33$
RSV-AZ-ST   2   5   4   4   9   1   6   6   6   1   1   1 $3.83$
RSV-FS-SN   9 13 12 13   7   5   1   1   1   7   7   7 $6.92$
RSV-FS-ST   5 11   9 10 11   3   2   2   2   6   3   5 $5.75$
RSV-GH-ST   1   3   3   3   5   2   7   7   7   2   2   6 $4.00$
EGARCH-N 13   9 13 12 16 17 11 11 11 10 18 17 $13.17$
EGARCH-T 12   8 11   8 10 12 14 17 15 13 16 16 $12.67$
REGARCH-N   8   6   7   7 18   9   9   9   9   9   9 10 $9.17$
REGARCH-T   7   4   6   6 17   8   8   8   8   8   4   8 $7.67$

### v4

@@ 4.3.2 VaR and ES forecasts
At the 5% level, the unconditional test shows that SV-GH-ST is significantly outperformed by SV-T, SV-AZ-SN, SV-AZ-ST, and all RSV and REGARCH models. In addition, RSV-GH-ST outperforms SV-N, and RSV-AZ-ST significantly outperforms RSV-AZ-SN. Among EGARCH and REGARCH models, the T specifications consistently outperform their N counterparts.
@@
For the conditional GW test at the 5% level, SV-GH-ST is significantly outperformed by RSV-AZ-SN, RSV-AZ-ST, RSV-FS-SN, RSV-FS-ST, and RSV-GH-ST. RSV-T performs significantly worse than RSV-GH-ST. EGARCH-N is significantly outperformed by EGARCH-T, and REGARCH-N by REGARCH-T.
@@
In summary, the results highlight the superior performance of RSV models—especially RSV-AZ-ST, RSV-FS-ST, and RSV-GH-ST—compared to SV, EGARCH, and REGARCH models, particularly under stringent risk levels. The inclusion of RV consistently improves the accuracy of tail risk forecasts. Moreover, as the FZ0 loss function jointly evaluates VaR and ES by emphasizing tail behavior, these results underscore the importance of capturing skewness and heavy tails in return distributions. The results for the N225 are broadly consistent with those for the DJIA at the 5% level and are therefore reported in the Supplementary Material.
@@
To explore differences among RSV models, we examine the distributional properties of their one-day-ahead return forecasts. Table 6 reports the mean and standard deviation of skewness, kurtosis, and several lower percentiles (0.1%, 1%, 5%, and 10%) based on the posterior predictive distributions. Figures 17 and 18 present the histograms of these statistics across the forecast horizon for the DJIA and N225, respectively.
@@
Remark.
Additional results for the COVID-19 period are reported in the Supplementary Material. While RSV models with skew-$t$ specifications improve forecast accuracy for volatility in the main sample, their relative performance for VaR and ES forecasts deteriorates during the COVID-19 period. This finding suggests that the benefits of modeling conditional asymmetry in the return distribution may be less pronounced under extreme market stress, where tail risk dynamics are dominated by abrupt and persistent shocks. A more detailed investigation of this issue is left for future research.
4.4 Summary
Table 7 presents the rankings of average losses for volatility, VaR, and ES forecasts. The average ranking in the final column highlights the overall superiority of the RSV models, with RSV-AZ-ST and RSV-GH-ST achieving the best performance.
Our findings demonstrate that RV is a valuable predictor for forecasting not only volatility but also tail risk measures such as VaR and ES. Overall, RSV models tend to deliver superior performance, particularly for the N225, while REGARCH-T remains highly competitive for the DJIA. Incorporating a skewed $t$-distribution within the RSV framework further enhances forecast accuracy, especially for tail risk measures.
@@ 4.4 Summary
SV-N 17 17 17 15 10 16 13 12 13 14 12 14 $14.17$
SV-T 14 14 14 14  7 10 18 17 16 15 15  9 $13.58$
SV-AZ-SN 18 18 18 18  1 13 14 14 14 16 13 15 $14.33$
SV-AZ-ST 15 15 15 16  4 12 17 16 17 17 14 13 $14.25$
SV-FS-SN 11  3  3  2  2 14 10 10 10 11 11 12 $8.25$
SV-FS-ST 10  1  2  1  3 15 12 13 12 12 10 11 $8.50$
SV-GH-ST 16 16 16 17 13 18 16 15 18 18 17 18 $16.50$
RSV-N  8 12 11 12 16  8  3  3  3  5  6  3 $7.50$
RSV-T  4  9  8  8 15  7  5  5  5  3  5  2 $6.33$
RSV-AZ-SN  5 10  9 10 14  4  4  4  4  4  8  4 $6.67$
RSV-AZ-ST  3  7  5  6 11  1  6  6  6  1  1  1 $4.50$
RSV-FS-SN  9 13 13 13  8  5  1  1  1  9  7  7 $7.25$
RSV-FS-ST  7 11 10 11 12  3  2  2  2  6  3  5 $6.17$
RSV-GH-ST  2  5  4  5  5  2  7  7  7  2  2  6 $4.50$
EGARCH-N 13  8 12  9 17 17 11 11 11 10 18 17 $12.83$
EGARCH-T 12  4  7  4  9 11 15 18 15 13 16 16 $11.67$
REGARCH-N  6  6  6  7 18  9  9  8  8  8  9 10 $8.67$
REGARCH-T  1  2  1  3  6  6  8  9  9  7  4  8 $5.33$